OpenStudy (anonymous):

solve for x : log2 3x=2=1 the first 2 is lower than the word log log5(2x+1)=1 5 is lower than the word log

OpenStudy (anonymous):

I don't know u write (3x+2) or (3x-2), If (3x+2) than x=0 \[\log _{2}^{(3x-2)}=1\]\[(3x-2)\log _{2}=\log(1)\]\[3x-2=(\log 1)(\log2)\]\[3x-2=2\]\[3x=2+2\]\[x=\frac{4}{3}\]and another\[\log _{5}(2x+1)=1\]\[(2x+1)\log _{5}=\log(1)\]\[2x+1=(\log(1)) \times(\log(5))\]\[2x+1=5\]\[2x=5-1\]\[x=\frac{4}{2}\]\[x=2\]

OpenStudy (anonymous):

lets try this. for the second one rewrite in equivalent exponential form. saying \[b^x=y\] same as \[\log_b(y)=x\]

OpenStudy (anonymous):

you have \[\log_5(2x+1)=1\] this means the same thing as \[2x+1=5^1=5\] and now you solve in your head to get \[x=2\]

OpenStudy (anonymous):

try as i might i cannot makes heads nor tails of the solution given above. first of all there is no such thing as \[log_5\] without something after it. second there is no way \[log_5(2x+1)=(2x+1)log_5\] this has not meaning at all. third \[log_5(1)=0\] because \[b^0=1\] and of course \[log_b(1)=0\] for all b.

OpenStudy (anonymous):

so this line \[2x+1=(log(1))×(log(5))\] says \[2x+1=0\]

OpenStudy (anonymous):

just remember to write in equivalent exponential form: \[log_b(x)=y \Leftrightarrow b^y=x\] and you should be able to switch back and forth with ease